OpenStudy (anonymous):

What is the inverse of f if f(x)=^3sqrtx-2?

OpenStudy (jdoe0001):

\(\large \bf f(x)=y=\sqrt[3]{x-2}\qquad \textit{inverse relationship}\implies x=\sqrt[3]{y-2}\) solve for "y"

OpenStudy (anonymous):

So would it be x=y^3+8

OpenStudy (jdoe0001):

well... x=y^3+8 <--- doesn't look solved for "y" for one.... and it wouldn't resolve to that

OpenStudy (anonymous):

I don't understand how to do it?